# Fine-grained Semantics Integration for Large Language Model-based Recommendation

### v1

@@ 1. Introduction
In a typical LLM-based generative recommendation pipeline (Liu et al., 2025; Kong et al., 2025b; You et al., 2025), the training procedure can be broadly divided into three stages: (1) item tokenization as semantic identifiers (SIDs)(Rajput et al., 2023; Wang et al., 2024; Hou et al., 2023): each item is first represented as a sequence of $N$ fixed-length tokens (i.e., SIDs) that lie outside the original vocabulary of the LLM (e.g., three tokens like ¡a_236¿ ¡b_231¿ ¡c_226¿), commonly obtained using vector quantization methods such as RQ-VAE(Zeghidour et al., 2021); (2) alignment(Zheng et al., 2024; Wang et al., 2024) between these out-of-vocabulary (OOV) item tokens and the LLM backbone through supervised fine-tuning (SFT), enabling the model to understand these tokens for recommendation; and (3) preference learning (e.g., reinforcement learning (RL)(Tan et al., 2025; Kong et al., 2025b; Team, 2025; Zhou et al., 2025; Kong et al., 2025a; You et al., 2025) or direct preference optimization (DPO) (Chen et al., 2024; Bai et al., 2024; Rafailov et al., 2023; Ye et al., 2025)) to further enhance the user preferences modeling capabilities, thereby pushing the upper bound of recommendation quality. As the first two steps, item tokenization and alignment with LLMs directly determine how the LLM understands items and thus largely affect recommendation performance, playing vital roles in LLM-based generative recommendation. However, existing approaches (Rajput et al., 2023; Zheng et al., 2024; Kong et al., 2025b) for these two steps remain coarse, which limits the LLM’s ability to effectively understand OOV item tokens and consequently leads to suboptimal recommendation performance. Specifically, we identify two key limitations in current item tokenization and alignment with LLMs as follows:
@@ 2.1. SID-based Generative Recommendation
2.2. LLM-based Generative Recommendation
@@ 3.2. Tokenization and Embedding Initialization
Let $\tilde{\mathbf{M}}=[\mathbf{x}_{1};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{N\times d}$ be the matrix stacking the embeddings of all $N$ items. We initialize $\mathbf{R}^{(1)}=\tilde{\mathbf{M}}$. For each layer $l\in\{1,\dots,L\}$, where $L$ is the number of hierarchical levels in the semantic codebook, we learn a codebook $\mathbf{C}^{(l)}=\{\mathbf{c}_{k}^{(l)}\}_{k=1}^{K_{l}}$ by running K-means with $K_{l}$ centroids on the current residuals $\mathbf{R}^{(l)}$:
@@
3.3. Supervised Fine Tuning for LLM-based Recommenders
@@ 4.1. SA-Init: Semantic-Aware embedding Initialization
    where $e_{i}$ represents the SID tuple of item $i$. This step groups items based on shared structural codes, serving as a proxy for shared semantics.

### v2

# arXiv:2602.22632v2[cs.IR] 28 Feb 2026

@@ 1. Introduction
In a typical LLM-based generative recommendation pipeline (Liu et al., 2025; Kong et al., 2025b; You et al., 2025), the training procedure can be broadly divided into three stages: (1) item tokenization as semantic identifiers (SIDs)(Rajput et al., 2023; Wang et al., 2024; Hou et al., 2023): each item is first represented as a sequence of NN fixed-length tokens (i.e., SIDs) that lie outside the original vocabulary of the LLM (e.g., three tokens like ¡a_236¿ ¡b_231¿ ¡c_226¿), commonly obtained using vector quantization methods such as RQ-VAE(Zeghidour et al., 2021); (2) alignment(Zheng et al., 2024; Wang et al., 2024) between these out-of-vocabulary (OOV) item tokens and the LLM backbone through supervised fine-tuning (SFT), enabling the model to understand these tokens for recommendation; and (3) preference learning (e.g., reinforcement learning (RL)(Tan et al., 2025; Kong et al., 2025b; Team, 2025; Zhou et al., 2025; Kong et al., 2025a; You et al., 2025) or direct preference optimization (DPO) (Chen et al., 2024; Bai et al., 2024; Rafailov et al., 2023; Ye et al., 2025)) to further enhance the user preferences modeling capabilities, thereby pushing the upper bound of recommendation quality. As the first two steps, item tokenization and alignment with LLMs directly determine how the LLM understands items and thus largely affect recommendation performance, playing vital roles in LLM-based generative recommendation. However, existing approaches (Rajput et al., 2023; Zheng et al., 2024; Kong et al., 2025b) for these two steps remain coarse, which limits the LLM’s ability to effectively understand OOV item tokens and consequently leads to suboptimal recommendation performance. Specifically, we identify two key limitations in current item tokenization and alignment with LLMs as follows:
@@ 2.1. SID-based Generative Recommendation
$2.$ LLM-based Generative Recommendation
@@ 3.2. Tokenization and Embedding Initialization
Let $\tilde{\mathbf{M}}=[\mathbf{x}_{1};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{N\times d}$ be the matrix stacking the embeddings of all NN items. We initialize $\mathbf{R}^{(1)}=\tilde{\mathbf{M}}$. For each layer $l\in\{1,\dots,L\}$, where $L$ is the number of hierarchical levels in the semantic codebook, we learn a codebook $\mathbf{C}^{(l)}=\{\mathbf{c}_{k}^{(l)}\}_{k=1}^{K_{l}}$ by running K-means with $K_{l}$ centroids on the current residuals $\mathbf{R}^{(l)}$:
@@
$3.$ Supervised Fine Tuning for LLM-based Recommenders
@@ 4.1. SA-Init: Semantic-Aware embedding Initialization
    where $e_{i}$ represents the SID tuple of item ii. This step groups items based on shared structural codes, serving as a proxy for shared semantics.